ON THE LIMIT BEHAVIOR OF THE SOLUTION OF A SEQUENCE OF GENERALIZED EIGENPROBLEMS IN DIFFERENT HILBERT

SPACES by Mladen Jurak, Zagreb, Croatia

Abstract. A sequence of generalized eigenproblems in different Hilbert spaces is considered and conditions ensuring convergence of eigenvalues and eigenvectors are given. An application to the homogenization problem is presented. Mathematics subject classification: 35P99, 35B27 Keywords: compact operator, Hilbert space, spectrum, homogenization.